Two numbers sum to 32. Six times the smaller number is equal to twice the larger. What are the numbers?
Answer by TutorDelphia(193)

Two numbers sum to 32
s+l=32
Six times the smaller number is equal to twice the larger.
6s=2l
divide both sides by 2
3s=l
sub in 3s for l
s+3s=32
combine like terms
4s=32
divide both sides by 4
s=8
sub in 8 for s
8+l=32
l=24